OpenStudy (gorica):

A particle mass m moves under the attractive force of intensity (ar+b^2)/r^3 per unit mass. How to show that this motion is planar? What does it mean "force of intensity (ar+b^2)/r^3 per unit mass"? The motion is in central field.

OpenStudy (anonymous):

torque about origin is zero as it is central motion. So, the direction as well as magnitude of angular momentum will remain constant. Now, \(\overrightarrow{L}=m\overrightarrow{r} \times\overrightarrow{v}\) Now\(\overrightarrow{r}\) is perpendicular to \(\overrightarrow{L}\) and \(\overrightarrow{L}\) is fixed in direction as i said above. This means \(\overrightarrow{r}\) is confined to rotate in the plane perpendicular to \(\overrightarrow{L}\).

OpenStudy (gorica):

Actually, I know how to show that motion in central field is planar. I thought that I have to use a given magnitude of force. Is there a way to show it for this motion under this force?

OpenStudy (anonymous):

Well, no matter what is the nature of a central force (attractive, repulsive, inverse squared or whatever), planar motion is its characteristic. Like u can see for this force. It can be written in this vector form. \(\huge{\overrightarrow{F}=-m\frac{(ar+b^{2})\hat{r}}{r^{3}}}\) \(\overrightarrow{\tau}=\overrightarrow{r}\times\overrightarrow{F}\) So, you can see that \(\overrightarrow{\tau}\) will always be zero and then the rest of the reasoning follows.

OpenStudy (anonymous):

Also your question is asking the meaning of "force per unit mass" which is clearly acceleration.

OpenStudy (gorica):

I noticed that, in problems, when they say "force of magnitude ______ per mass unit", in Binet's equation there's no m.